If a firm manager has a base salary of $100,000 and also receives 6 percent of all profits, what percentage of his/her final income will be from a profit-sharing plan when profit equals $1,500,000?


1
Expert's answer
2021-11-22T10:04:22-0500

Basic Salary= $100,000.

Extra Earnings = 6% of $1,500,000.

"=0.06\times 1,500,000=90,000."

Final Income"=100,000+90,000"

"=190,000"

The percentage of final income from the profit- sharing plan:

"=\frac{190,000}{1,500,000}\times100=12.67"%